3-2 (2x-1) is great than or equal to 3(2x-2) +1
Answer by edjones(8007)

3-2(2x-1)>=3(2x-2)+1
3-4x+2>=6x-6+1
-4x+5>=6x-5
-10x>=-10
x<=1 Don't forget to reverse the inequality sign when dividing by a negative number.
.
Ed